Hi, my question is this
Two aircraft both depart from the same airport at different times. The 1st takes off at 14:00 (2pm) and maintains a speed of 180km/h. The second takes off at 19:00 (7pm) and travels at 480km/h. By which time will the second aircraft catch up with the first?
Thought I was a whizz at Speed/Distance/Time calculation but this one took me by suprise. The answer in Laymans terms and Algebra would be vastly appreciated!

Answer by htmentor(1343)

Two aircraft both depart from the same airport at different times. The 1st takes off at 14:00 (2pm) and maintains a speed of 180km/h. The second takes off at 19:00 (7pm) and travels at 480km/h. By which time will the second aircraft catch up with the first?
======================================================
The 1st plane has a head start of 180 km/h * 5 h = 900 km
The RELATIVE speed of the two planes is 480 km/h - 180 km/h = 300 km/h
This is as if plane 1 were standing still and plane 2 approached it at 300 km/h.
So the time required to travel the 900 km is:
t = d/v = 900 km/300 km/h = 3 h
Therefore, the time will be 7 pm + 3 h = 10 pm